OpenStudy (kimie):

A spinner has three congruent sections numbered 1, 2, 3. Yolanda made 50 spins of this spinner and recorded the outcomes in this table. Outcome Number 1 15 2 19 3 16 What is the theoretical probability of the spinner stopping on 2? What is the experimental probability of the spinner stopping on 2? A. Theoretical - 38% Experimental - 33.3% B. Theoretical - 33.3% Experimental - 19% C. Theoretical - 16.7% Experimental - 19% D. Theoretical - 33.3% Experimental - 38%

OpenStudy (kimie):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

`What is the theoretical probability of the spinner stopping on 2? ` there are \(\LARGE \color{green}{3}\) sections total. In this case, \(\LARGE \color{red}{1}\) of them is labeled "2" divide the values \[\Large \frac{\color{red}{\text{number of sections labeled '2'}}}{\color{green}{\text{number of sections total}}}=\frac{\color{red}{1}}{\color{green}{3}} \approx 0.333 \approx 33.3\%\]

jimthompson5910 (jim_thompson5910):

For the experimental probability, you'll divide (number of times a "2" comes up)/(number of spins total)

OpenStudy (kimie):

Woow

OpenStudy (kimie):

okay, so 19/50?

jimthompson5910 (jim_thompson5910):

yep 19/50 = 0.38 = 38%

OpenStudy (kimie):

Oh! Okay! Thank you!

jimthompson5910 (jim_thompson5910):

or you can think of it as 19/50 = 38/100 ... multiply top and bottom by 2 38/100 = 38%

jimthompson5910 (jim_thompson5910):

you're welcome